# Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy

### v1

@@ Abstract
This paper identifies that a group of state-of-the-art locally-differentially-private (LDP) algorithms for frequency estimation are equivalent to the private Count-Mean Sketch (CMS) algorithm with different parameters. Therefore, we revisit the private CMS, correct errors in the original CMS paper regarding expectation and variance, modify the CMS implementation to eliminate existing bias, and explore optimized parameters for CMS to achieve optimality in reducing the worst-case mean squared error (MSE), $l_{1}$ loss, and $l_{2}$ loss. Additionally, we prove that pairwise-independent hashing is sufficient for CMS, reducing its communication cost to the logarithm of the cardinality of all possible values (i.e., a dictionary). As a result, the aforementioned optimized CMS is proven theoretically and empirically to be the only algorithm optimized for reducing the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss when dealing with a very large dictionary. Furthermore, we demonstrate that randomness is necessary to ensure the correctness of CMS, and the communication cost of CMS, though low, is unavoidable despite the randomness being public or private.
@@ 1 Introduction
One of the most popular approaches to protect data is differential privacy [1], which is considered the de facto standard for extracting statistics from a dataset while protecting individual data privacy. Differential privacy has a local version [2], known as Local Differential Privacy (LDP), which requires enough perturbation on individual data such that an analyst or attacker cannot gain significant knowledge about the individual data values despite their prior knowledge [3]. However, LDP also adds more challenges to calculating statistics from a dataset, including frequency estimation.
@@
A major challenge of LDP-preserving frequency estimation is that traditional mechanisms, such as randomized response [4], result in the variance of the estimated frequency increasing linearly with the number of all possible values in a dataset (referred to as a dictionary), making the estimation infeasible if the dictionary is too large. Several algorithms [5, 6, 7, 8, 9, 10, 11, 12, 13] have been developed to address this problem, which will be discussed in the Previous Work (Section 1.1). Many of these algorithms (e.g., [5, 9, 14, 12]) are found to be equivalent to the Count-Mean Sketch (CMS) algorithm (see Section 4 for more details). CMS hashes the original values through randomly selected hash functions and estimates the original frequencies based on the frequencies of the hashed values. The LDP version of CMS was originally developed by [7], but it had minor issues in calculating the expectation and variance (see Section 3.6) and did not explore the optimized parameters for CMS. Therefore, our work provides a corrected calculation of expectation and variance (Sections 3.1 and 3.2) and explores the optimized parameters of CMS (Section 3.4), which optimally reduce the worst-case mean squared error ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss as follows:
$\widehat{MSE}^{*}=\begin{cases}\frac{2(\Delta_{MSE}+e^{\epsilon})}{n(e^{\epsilon}-1)^{2}}&\text{ if }f^{*}\leq\frac{1}{2}\\ \frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}&\text{ if }\frac{1}{2}<f^{*}\leq 1\end{cases}$
$l_{1}^{*}\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}}\xrightarrow{d\rightarrow\infty}\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)}$
$l_{2}^{*}=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}\xrightarrow{d\rightarrow\infty}\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
with
$\Delta_{MSE}=e^{\epsilon/2}\sqrt{[(1-f^{*})e^{\epsilon}+f^{*}][f^{*}e^{\epsilon}+(1-f^{*})]}$
$\Delta_{l}=e^{\epsilon/2}\sqrt{(e^{\epsilon}+d-1)(de^{\epsilon}-e^{\epsilon}+1)}$
@@
where $\epsilon$, $n$, and $d$ are the privacy factor, dataset size, and dictionary size, respectively; $f^{*}$ is defined as $\forall x\in\mathbf{x}:f(x)\leq f^{*}$ where $\mathbf{x}$ are the values of interest; * denotes the optimality. Moreover, our work proves that pairwise-independent hash functions are sufficient for CMS, and Section 3.5 proposes a practical approach to construct a pairwise-independent hashing family given any range of a hash function. The pairwise independence can reduce the communication cost to the logarithm of the dictionary size. As a result, Section 5.1 compares various state-of-the-art algorithms and concludes that our optimized CMS is the only algorithm optimized in the worst-case MSE, $l_{1}$ loss, and $l_{2}$ loss while its communication cost is only logarithmic relative to the dictionary size. Additionally, Section 5.2 proves the necessity of randomness in CMS and provides a counterexample against the existing deterministic algorithm. Section 5.3 shows that the communication cost of CMS is unavoidable, despite the randomness being public or private. Section 6 empirically demonstrates that the optimized CMS outperforms other algorithms.
@@ 1.1 Previous Work
RAPPOR [6] is the first algorithm to estimate frequency with accuracy independent of dictionary size. It represents the original value as a one-hot vector and applies LDP perturbation to every bit of the array. However, the communication cost is proportional to the dictionary size, as the one-hot vector is the same size as the dictionary. [15] developed another version of RAPPOR, which hashes an original value to a smaller range than the dictionary size, such that the one-hot encoding results in a much smaller vector. This version uses gradient descent to search for one possible solution for the original frequencies. However, this mapping is equivalent to dimension reduction in linear algebra, so the found solution is not guaranteed to be close to the original frequencies. To tackle the large dictionary, [5] developed an algorithm that transforms an original value to a column of a Hadamard matrix, uniformly samples one element from the column, and perturbs it. The original frequencies can be reconstructed from the transformed values by applying the Hadamard matrix again. The accuracy of this approach is independent of the dictionary size, and the communication cost is the logarithm of the dictionary size. The Apple Privacy team [7] modified the Count-mean Sketch (CMS) algorithm to be LDP. However, they failed to derive the expectation and variance correctly (see Section 3.6) and did not explore the optimized settings of CMS. In addition to CMS, other members of the Count Sketch family have also been adapted to LDP (e.g., [8, 13]). [9] summarized existing approaches and proposed a variant of RAPPOR and a local hashing method, which were claimed to be optimal in terms of MSE. However, [9] assumed the estimated frequencies to be always close to zero, so their optimality did not hold when the original frequency approaches one. At the same time, [10] proposed a Subset Selection algorithm to estimate frequency with optimized accuracy in $l_{1}$ / $l_{2}$ losses, but its communication cost is also proportional to the dictionary size when the privacy requirement is high. [11] later integrated the Hadamard matrix into the subset selection and reduced the communication cost to the logarithm of the dictionary size. [12] further improved this method by applying the Hadamard transform twice. Though its performance can achieve order-optimal in $l_{1}$ / $l_{2}$ losses, it may perform worse than the Hadamard Response in terms of MSE (as discussed in Section 4.3).
@@ 2.1 Local Differential Privacy (LDP)
Local Differential Privacy (LDP) was initially proposed in [2] to ensure that querying an object does not significantly disclose its actual value.
@@
A mechanism $M$ is $\epsilon$-LDP if and only if $\forall(v,v^{\prime})\in\mathcal{V}^{2}$, and $\forall u\in\mathcal{U}$,
$Pr(M(v)=u|v)\leq e^{\epsilon}Pr(M(v^{\prime})=u|v^{\prime}).$
While there are other variations of LDP, this paper focuses solely on the most widely-used form mentioned above.
Given the above definition, $P$ is defined as a $|\mathcal{U}|\times|\mathcal{V}|$ matrix, where $P[u,v]$ represents the probability of the LDP mechanism $M$ generating $u$ given the input $v$. Here $u$ and $v$ are represented by an index between $1..|\mathcal{U}|$ and $1..|\mathcal{V}|$, respectively. Suppose there exists another matrix $Q$ of shape $|\mathcal{V}|\times|\mathcal{U}|$, which satisfies $I=QP$ and $I$ is an identity matrix. The decoding process of $M$ is defined to output $Q[:,u]$ given the output $u$ from $M$. The reconstruction process (perturbation and decoding) is denoted as $R$, whose output is an estimator of the input $v$, formulated as:
@@
$Pr(M(v)=u|v)=Pr(M(v)=u|v,\theta),$
where $\theta$ represents any other information (except the randomness generator used by $M$). Therefore, the outputs of $M$ from two inputs are mutually independent, which is formulated as:
@@
If $v$ and $v^{\prime}$ are two inputs into $M$, the reconstruction of them from the outputs of $M$ satisfies $\forall(a,b)\in\mathcal{V}^{2}:Cov(R(v)[a]),R(v^{\prime})[b])=0$.
Most widely-used algorithms (e.g., randomized response [4] and RAPPOR [6]) satisfy this property. Additionally, we observe that these LDP mechanisms also exhibit symmetry in variance, formulated as:
@@
When $\forall v,v^{\prime}\in\mathcal{V}^{2}$ and $v\neq v^{\prime}$, $Var(R(v)[v^{\prime}])$ are identical, denoted as $Var(R|\neq)$; $\forall v\in\mathcal{V}$, $Var(R(v)[v])$ are also identical, denoted as $Var(R|=)$.
@@ 2.2 Frequency Estimation
Let’s define some symbols: $X^{(i)}$ is the $i$-th object of an arbitrary dataset of size $n$; the possible values of all the objects comprise an dictionary $[d]$, where $d$ is its size, and $\forall i\in[n]:X^{(i)}\in[d]$. The frequency of value $x$ is defined as $f(x)=\frac{\sum_{i\in[n]}\mathbf{1}\{X^{(i)}=x\}}{n}$. The LDP workflow of frequency estimation usually proceeds as follows:
1. Each object $X^{(i)}$ is perturbed by an LDP mechanism, which outputs $Y^{(i)}$.
2. The server collects and aggregates the outputs as $\hat{f}(x)=A_{x}([Y^{(i)}]_{i\in[n]})$, if the server is interested in $f(x)$. The aggregation can be run for every $x\in[d]$.
$\hat{f}(x)$ is an estimator of $f(x)$. Existing literature usually uses $l_{1}$ and $l_{2}$ losses to measure its overall accuracy, which are defined as follows:
@@
where $\mathbf{X}=[X^{(i)}]_{i\in[n]}$, and $\forall\mathbf{X}$ assumes the size $n$ is constant. These two losses measure the overall accuracy of the estimation. However, in many cases, an analyst is also interested in the accuracy of the frequency estimation of an individual value. The mean squared error (MSE) is commonly used to measure the accuracy of an estimation, reflecting both the variance and bias of the estimation. Since an analyst is assumed to have no knowledge of the actual dataset, they need to assume the worst case when calculating the MSE, which is defined as
$\widehat{MSE}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\max_{x\in\mathbf{x}}E[(\hat{f}(x)-f(x))^{2}],$ (1)
where $\widehat{MSE}$ denotes the upper bound of MSE, i.e., the worst case; $\mathbf{x}$ denotes the values of interest, which is $[d]$ by default. If the estimator $\hat{f}$ is unbiased, the worse-case MSE is the same as the worst-case variance, which is also positively correlated to the confidence interval (CI). Thus, an unbiased estimator optimized for $\widehat{MSE}$ is also optimized for CI.
@@ 3 Revisit Count-mean Sketch
3. Collect all the responses $(z^{(i)})_{i\in[n]}$, decode them, and acquire the corresponding $\hat{y}^{(i)}$ as an estimator of $y^{(i)}$, i.e., $\hat{y}^{(i)}=R(y^{(i)})$, where $R$ is defined in Section 2.1. Notice that each $\hat{y}^{(i)}$ is a vector of size $m$.

### v2

# Improving Count-Mean Sketch as the Leading Locally Differentially Private Frequency Estimator for Large Dictionaries

@@ Abstract
This paper identifies that a group of latest locally-differentially-private (LDP) algorithms for frequency estimation, including all the Hadamard-matrix-based algorithms, are equivalent to the private Count-Mean Sketch (CMS) algorithm with different parameters. Therefore, we revisit the private CMS, correct errors in the original CMS paper regarding expectation and variance, modify the CMS implementation to eliminate existing bias, and optimize CMS using randomized response (RR) as the perturbation method. The optimized CMS with RR is shown to outperform CMS variants with other known perturbations in reducing the worst-case mean squared error (MSE), $l_{1}$ loss, and $l_{2}$ loss. Additionally, we prove that pairwise-independent hashing is sufficient for CMS, reducing its communication cost to the logarithm of the cardinality of all possible values (i.e., a dictionary). As a result, the optimized CMS with RR is proven theoretically and empirically as the leading algorithm for reducing the aforementioned loss functions when dealing with a very large dictionary. Furthermore, we demonstrate that randomness is necessary to ensure the correctness of CMS, and the communication cost of CMS, though low, is unavoidable despite the randomness being public or private.
@@ 1 Introduction
One of the most popular approaches to protect data is differential privacy (DP) [1], which is considered the de facto standard for extracting statistics from a dataset while preserving individual data privacy. DP introduces sufficient noise to the statistical outputs such that the participation of any individual’s data cannot be confidently inferred from the query results. However, DP still requires individuals to upload their data to a server before extracting statistics, posing risks of eavesdropping during data transmission or data breaches on the server side [2]. To address these concerns, Local Differential Privacy (LDP) [3] has been proposed, which requires enough perturbation on individual data on the client side before sending them to the server for statistical analysis, theoretically preventing eavesdropping or data breaches. However, LDP also adds more challenges to calculating statistics from a dataset, including frequency estimation.
@@
A major challenge of LDP-preserving frequency estimation is that traditional mechanisms, such as randomized response [4], result in the precision of the estimated frequency increasing linearly with the number of all possible values in a dataset (referred to as a dictionary), making the estimation infeasible if the dictionary is too large. Several algorithms [5, 6, 7, 8, 9, 10, 11, 12, 13] have been developed to address this problem, and their precision is proven to be independent of the dictionary size, though some of them still have communication cost linear to the dictionary size. These algorithms will be further introduced in the Previous Work (Section 1.1). Many of these algorithms (e.g., [5, 9, 14, 12]) are found to be equivalent to the Count-Mean Sketch (CMS) algorithm (see Section 4 for more details). CMS hashes the original values through randomly selected hash functions and estimates the original frequencies based on the frequencies of the hashed values. The LDP version of CMS was originally developed by [7], but it had minor issues in calculating the expectation and variance (see Section 3.7) and did not explore the optimized parameters for CMS. Therefore, our work provides a corrected calculation of expectation and variance (Sections 3.1 and 3.2) and optimize CMS using randomized response as the perturbation method (referred to as OCMS+RR). OCMS+RR is shown to outperform CMS variants with other known perturbations in reducing the worst-case mean squared error ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss (Section 3.4). Moreover, our work proves that pairwise-independent hash functions are sufficient for CMS, and Section 3.5 proposes a practical approach to construct a pairwise-independent hashing family given any range of a hash function. The pairwise independence can reduce the communication cost to the logarithm of the dictionary size, allowing OCMS+RR to be applied to datasets with very large dictionaries. The pseudocode of OCMS+RR is presented in Section 3.6 (source code available at https://github.com/mingen-pan/CountMeanSketch). When the dictionary size is much larger than the privacy factor $\epsilon$ defined by LDP, the aforementioned loss functions (defined in Section 2.2) are reduced by OCMS+RR as
$\widehat{MSE}\leq\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$
$l_{1}\xrightarrow{d\gg e^{\epsilon}}\frac{2de^{\epsilon/2}}{\sqrt{n}(e^{\epsilon}-1)}$
$l_{2}\xrightarrow{d\gg e^{\epsilon}}\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$
@@
where $n$, and $d$ are the dataset size and dictionary size, respectively. As a result, Section 5.1 compares existing algorithms with OCMS+RR and concludes that our OCMS+RR is the leading algorithm for reducing the worst-case MSE ($\widehat{MSE}$), $l_{1}$ loss, and $l_{2}$ loss, while its communication cost is only logarithmic relative to the dictionary size. Additionally, Section 5.2 proves the necessity of randomness in CMS and provides a counterexample to an attempt to de-randomize a CMequivalent algorithm by [14]. Section 5.3 shows that the communication cost of CMS is unavoidable, despite the randomness being public or private. Section 6 empirically demonstrates that OCMS+RR outperforms other algorithms.
@@ 1.1 Previous Work
RAPPOR [6] is the first algorithm to estimate frequency with precision independent of dictionary size. It represents the original value as a one-hot vector and applies LDP perturbation to every bit of the array. However, the communication cost is proportional to the dictionary size, as the one-hot vector is the same size as the dictionary. An alternative version of RAPPOR [15] hashes an original value into a smaller range than the dictionary size, resulting in a much smaller one-hot encoded vector. This version uses gradient descent to search for one possible solution for the original frequencies. However, this mapping is equivalent to dimension reduction in linear algebra, so the found solution is not guaranteed to be close to the original frequencies. To tackle the large dictionary, an algorithm was developed by [5], which transforms an original value to a column of a Hadamard matrix, uniformly samples one element from the column, and perturbs it. The original frequencies can be reconstructed from the transformed values by applying the Hadamard matrix again. The precision of this approach is independent of the dictionary size, and the communication cost is the logarithm of the dictionary size. The Apple Privacy team [7] modified the Count-mean Sketch (CMS) algorithm to be LDP. However, they failed to derive the expectation and variance correctly (see Section 3.7) and did not explore the optimized settings of CMS. In addition to CMS, other members of the Count Sketch family have also been adapted to LDP (e.g., [8, 13]). Wang et al. [9] summarized existing approaches and proposed a variant of RAPPOR and a local hashing method, which were claimed to be optimal in terms of MSE. However, [9] assumed the estimated frequencies to be always close to zero, so their optimality did not hold when the original frequency approaches one. Simultaneously, the Subset Selection algorithm [10] was proposed to estimate frequency with leading precision in $l_{1}$ / $l_{2}$ losses, but its communication cost is also proportional to the dictionary size when the privacy requirement is high. Acharya et al. [11] later integrated the Hadamard matrix into the subset selection and reduced the communication cost to the logarithm of the dictionary size. This method was further improved by [12], which applies the Hadamard transform twice. Though its performance can achieve the state-of-the-art order of precision in $l_{1}$ / $l_{2}$ losses, it may perform worse than the Hadamard Response in terms of MSE (as discussed in Section 4.3).
@@ 2.1 Local Differential Privacy (LDP)
Local Differential Privacy (LDP) was initially proposed in [3] to ensure that querying an object does not significantly disclose its actual value. Consider a variable $V$ that is perturbed by a mechanism $M$:
@@
$M$ is $\epsilon$-LDP if and only if $\forall v,v^{\prime}\in\mathcal{V}$, and $\forall u\in\mathcal{U}$,
$Pr(M(V)=u|V=v)\leq e^{\epsilon}Pr(M(V)=u|V=v^{\prime}),$
where $\mathcal{V}$ and $\mathcal{U}$ represent the dictionary (all possible values) of $V$ and the range of $M$, respectively. This paper may abbreviate $\left(M\left(V\right)|V=v\right)$ as $M(v)$ when the variable is not the focus and its value is known.
Given the above definition, $P$ is defined as a $|\mathcal{U}|\times|\mathcal{V}|$ matrix, where $P[u,v]=Pr(M(v)=u)$. Here $u$ and $v$ are represented by an index between $1..|\mathcal{U}|$ and $1..|\mathcal{V}|$, respectively. We consider only cases where $|\mathcal{U}|\geq|\mathcal{V}|$ and $P$ is full rank, i.e., $\text{rank}(P)=|\mathcal{V}|$. This condition should hold for all meaningful LDP mechanisms, as the expected frequencies of the perturbed values correspond to a linear transformation of the original frequencies, and a rank-deficient transformation would prevent the reconstruction of the original frequencies from the perturbed values. Widely-used algorithms (e.g., [4, 6]) satisfy this condition. Subsequently, a matrix $Q$ is constructed as $(P^{T}P)^{-1}P^{T}$, which satisfies $I=QP$ with $I$ being an identity matrix. Given an output $u$ from $M$, its decoding is defined to output $Q[:,u]$, i.e., the $u$-th column of $Q$. Thus, the reconstruction process (perturbation and decoding) of a variable $V$ is defined as $R(V)=Q[:,M(V)]$. Similarly, $\left(R\left(V\right)|V=v\right)$ is abbreviated as $R(v)$ when the variable is not the focus and its value is known. An example of the reconstruction is presented in Appendix A. The randomness of $R(v)$ comes from $M(v)$, and each possible value of $R(v)$ is a vector of size $|\mathcal{V}|$ satisfying:
@@
$Pr(M(V)=u|V=v)=Pr(M(V)=u|V=v,\theta),$
where $\theta$ represents any other information (except the randomness generator used by $M$). Considering two variables $V$ and $V^{\prime}$, whose values are determined prior to perturbation, the stateless requirement ensures that their reconstruction process $R(V)$ and $R(V^{\prime})$ are mutually independent random variables, formulated as
@@
The covariance of $R(V)$ and $R(V^{\prime})$ satisfies $\forall v,v^{\prime},v_{a},v_{b}\in\mathcal{V}:Cov(R(V)[v_{a}],R(V^{\prime})[v_{b}]|V=v,V^{\prime}=v^{\prime})=0$.
Widely-used algorithms (e.g., [4, 6]) satisfy this property. Additionally, we observe that these LDP mechanisms also exhibit symmetry in variance, denoted as $Var$, formulated as:
@@
Every element of $\{Var(R(v)[v^{\prime}])|v,v^{\prime}\in\mathcal{V},v\neq v^{\prime}\}$ has an identical value, denoted as $Var(R|\neq)$; every element of $\{Var(R(v)[v])|v\in\mathcal{V}\}$ has an identical value, denoted as $Var(R|=)$.
@@ 2.2 Frequency Estimation
Let’s define some symbols: $X^{(i)}$ is the $i$-th object of an arbitrary dataset of size $n$; the possible values of all the objects comprise a dictionary $[d]$, where $d$ is its size, and $\forall i\in[n]:X^{(i)}\in[d]$. The frequency of value $x$ is defined as $f(x)=\frac{\sum_{i\in[n]}\mathbf{1}\{X^{(i)}=x\}}{n}$. The LDP workflow of frequency estimation typically proceeds as follows:
1. Each object $X^{(i)}$ is perturbed by an LDP mechanism, which outputs $\tilde{X}^{(i)}$.
2. An analyst collects and aggregates the outputs as $\hat{f}(x)=A_{x}([\tilde{X}^{(i)}]_{i\in[n]})$, if the analyst is interested in $f(x)$, where $A_{x}$ denotes the aggregation. The aggregation can be performed for every $x\in[d]$.
A potential scenario involves a group of clients sending their perturbed responses to a server, while their unperturbed responses constitute a dataset. The server then estimates the frequency of the unperturbed responses.
$\hat{f}(x)$ in Step 2 is an estimator of $f(x)$. Existing literature usually uses $l_{1}$ and $l_{2}$ losses to measure its overall precision, which are defined as follows:
@@
where $\mathbf{X}=[X^{(i)}]_{i\in[n]}$, and $\forall\mathbf{X}$ assumes the size $n$ is constant. These two losses measure the overall precision of the estimation. However, in many cases, an analyst is also interested in the precision of the frequency estimation of an individual value. The mean squared error (MSE), formulated as $MSE(\hat{f}(x))=E[(\hat{f}(x)-f(x))^{2}]$, is commonly used to measure the precision of an estimation, reflecting both the variance and bias of the estimation. Since an analyst is assumed to have no knowledge of the actual dataset, they need to assume the worst case when calculating the MSE, which is defined as
$\widehat{MSE}(\hat{f})=\max_{\forall\mathbf{X}:|\mathbf{X}|=n}\max_{x\in\mathbf{x}}MSE(\hat{f}(x)),$ (1)
where $\widehat{MSE}$ denotes the upper bound of MSE, i.e., the worst case; $\mathbf{x}$ denotes the values of interest, which is $[d]$ by default. If the estimator $\hat{f}$ is unbiased, the worst-case MSE is the same as the worst-case variance, which is also positively correlated to the confidence interval (CI). Thus, an unbiased estimator optimized for $\widehat{MSE}$ is also optimized for CI.
@@ 3 Revisit Count-mean Sketch
3. Collect all the responses $(z^{(i)})_{i\in[n]}$, decode them following Section 2.1 as $(\hat{y}^{(i)})_{i\in[n]}$, where each $\hat{y}^{(i)}$ is an estimator of $y^{(i)}$, i.e., $\hat{y}^{(i)}=R(y^{(i)})$, with $R$ defined in Section 2.1. Notice that each $\hat{y}^{(i)}$ is a vector of size $m$.